# Quantum algorithms in distributed quantum computing

### v2

@@ 4.3 Distributed quantum amplitude estimation
where $a$ is the unknown amplitude to be estimated, and $|\psi_{0}\rangle_{n}$ and $|\psi_{1}\rangle_{n}$ are (normalized) good and bad states, respectively, that correspond to the probability of measurement of 00$\rangle$0 ⟩ and $|1\rangle$ in the last qubit, respectively.
QAE has applications such as calculating the expectation value of the discrete linear function, quantum Monte Carlo integration [Nik2020, Herbert2022], quantum risk analysis, and portfolio optimization. We use the DQCS to simulate distributed QAE (DQAE) using multiple nodes. There are multiple approaches to performing QAE. Here, we study DQAE using maximum likelihood estimation [Suzuki2020], which achieves quadratic speedup over classical Monte Carlo algorithms in DQC without using the QFT. We show that the estimation error for the DQAE [Suzuki2020], simulated using the DQCS, achieves the Cramer-Rao bound, which provides a way to scale QAE using multiple QPUs.
We prepare the $A$ operator, using the probability distribution loading $P(x)$ [Nik2020], that loads a discrete probability distribution to a quantum state. The bounded function $f(x)$ is constructed from the samples of $P(x)$ and realized using multiple controlled-$R_{y}$ gates with controlled-$R_{y}(f(i))$ on the gate on the $ith$ qubit [Nik2020]. The probability of the measurement of $|1\rangle$ on the last qubit gives the expectation value $\mathrm{E}[f(x)]$.
@@
Suppose P(x) is a uniform distribution, the probability of measuring $|1\rangle$ on the last qubit is given by $\frac{1}{2^{n}}\sum_{i=0}^{2^{n}-1}f(x_{i})$ for $x_{i}=i/2^{n}$ [Carrera2021]. For QAE, we construct the $Q$ operator, such that $Q=-AS_{0}A^{-1}S_{X}$. The operator $S_{X}$ leaves the good states unchanged and includes a phase $(-1)$ on the bad states. $S_{X}|\psi_{0}\rangle|0\rangle=|\psi_{0}\rangle|0\rangle$, $S_{X}|\psi_{1}\rangle|1\rangle=-|\psi_{1}\rangle|1\rangle$. The operator $S_{0}$ attaches a phase (-1) to $|0\rangle_{n}$ and leaves other basis states unchanged, the repeated application of $Q$, $M$ times, on the state $A|0\rangle$ gives
@@
QAE requires multiple controlled unitary gates for the implementation of the reflection operators. We could write multiple CNOT gates (with $n$ control qubits and a single-qubit unitary) with single-qubit gates and $\leq(20n-38)$ CNOT gates [Rv2023, Eisert2000]. We consider the QAE without phase estimation that uses the maximum likelihood estimate [Suzuki2020] for the integration of $\int_{0}^{c}\sin^{2}(x)$ in DQC. We compare the estimated error values against the Cramer-Rao bound [Suzuki2020], where $P(x)$ is a uniform distribution obtained by applying Hadamard gates to all of the qubits. The function $f(x)$ is prepared with $a=c/2^{n-1},b=c/2^{n}$ using 1 $R_{y}$ gate and $n$ controlled-Ry gates, $Ry(c/2^{n-i})$ [Suzuki2020]. It can be seen in Fig. 7 (($c=\frac{\pi}{3}$)) that DQAE can be simulated and the estimated value achieves the Cramer-Rao bound.
@@
In our simulation, we used Qiskit’s transpile function to apply a distributed multiple controlled gate and distributed controlled unitary gates using non-local CNOT gates (Fig. 2). The DQCS can be used to obtain the resource estimation for the DQAE for various node parameters in Table 1.
@@
Table 1: Resource estimation for the distributed quantum amplitude estimation (DQAE) obtained using the DQCS where the number of nodes $(n)$, number of Grover operators $(m)$, number of entanglement generation steps $k$ for the qubit-node coupling parameter $p=(1,0.8,0.5)$, and estimation error for $m=1$
are considered
@@
loaded to quantum state $|p\rangle=\sum_{i=0}^{2^{d-1}}\sqrt{p_{i}}|x_{i}\rangle$, where $p_{i}$ is sampled from $P(x)$ with $2^{d}$ points [Nik2020, Carrera2021]. We use the DQCS to simulate the quantum state preparation of a probability distribution. We consider a normal distribution with $8$ qubits with the mean $\mu=1$ and $\sigma=2$ in the DQCS, with $n=1,2,4,8$ nodes, to obtain the Heilinger fidelity in Fig. 8.
Figure 8: The Heilinger fidelity of a normal distribution with 8 qubits with different numbers of nodes, where $\epsilon=(0.002(\text{blue}),0.005(\text{orange}),0.009(\text{green}))$
@@ 5 Conclusion
We outlined a distributed quantum circuit simulator (DQCS), which uses a number of interconnected quantum processing units (QPUs) consisting of local and communication qubits. We studied foundational quantum algorithms such as the quantum Fourier transform, quantum phase estimation, and amplitude estimation using the DQCS and showed the application of dynamic quantum circuits in DQC. Recently, there has been growing interest in the realization of quantum networks and distributed quantum computing in various platforms such as superconducting qubits, NV centers [Nature2013, NV], ions [PhysRevLett.124.110501], and neutral atoms [neutral]. The DQCS provides a platform for users to input the experiment parameters, simulate distributed quantum circuits, and obtain the estimates for quantum algorithms including quantum phase estimation, quantum amplitude estimation, quantum state preparation of the probability distribution, etc. In the future, we aim to integrate quantum internet simulators with the DQCS to study distributed quantum algorithms in the quantum internet. We look forward to including quantum error correction and fault tolerance in the DQCS to obtain favourable polynomial scaling for the fidelity of quantum algorithms [Muralidharan2016]. We could also study distributed quantum machine learning algorithms for classification and regression [Tang2023].
@@ Acknowledgements
I would like to thank Constantin Gonciulea, Charlee Stefanski, Vanio Markov and David Novak for helpful discussions, comments on the DQCS, and suggestions to improve the quality of the manuscript.

### v3

@@ 4.3 Distributed quantum amplitude estimation
where $a$ is the unknown amplitude to be estimated, and $|\psi_{0}\rangle_{n}$ and $|\psi_{1}\rangle_{n}$ are (normalized) good and bad states, respectively, that correspond to the probability of measurement of $|0\rangle$ and $|1\rangle$ in the last qubit, respectively.
QAE has applications such as calculating the expectation value of the discrete linear function, quantum Monte Carlo integration [Nik2020, Herbert2022], quantum risk analysis, and portfolio optimization. Using multiple nodes, we use the DQCS to simulate distributed QAE (DQAE). There are various approaches to performing QAE. Here, we study DQAE using maximum likelihood estimation [Suzuki2020], which achieves quadratic speedup over classical Monte Carlo algorithms in DQC without using the QFT. We show that the estimation error for the DQAE [Suzuki2020], simulated using the DQCS, achieves the Cramer-Rao bound, which provides a way to scale QAE using multiple QPUs.
We prepare the $A$ operator, which loads a discrete probability distribution to a quantum state using the probability distribution loading $P(x)$ [Nik2020]. The bounded function $f(x)$ is constructed from the samples of $P(x)$ and realized using multiple controlled-$R_{y}$ gates with controlled-$R_{y}(f(i))$ on the gate on the $ith$ qubit [Nik2020]. The probability of the measurement of $|1\rangle$ on the last qubit gives the expectation value $\mathrm{E}[f(x)]$.
@@
Suppose P(x) is a uniform distribution, the probability of measuring $|1\rangle$ on the last qubit is given by $\frac{1}{2^{n}}\sum_{i=0}^{2^{n}-1}f(x_{i})$ for $x_{i}=i/2^{n}$ [Carrera2021]. For QAE, we construct the $Q$ operator, such that $Q=-AS_{0}A^{-1}S_{X}$. The operator $S_{X}$ leaves the good states unchanged and includes a phase $(-1)$ on the bad states. $S_{X}|\psi_{0}\rangle|0\rangle=|\psi_{0}\rangle|0\rangle$, $S_{X}|\psi_{1}\rangle|1\rangle=-|\psi_{1}\rangle|1\rangle$. The operator $S_{0}$ attaches a phase (-1) to $|0\rangle_{n}$ and leaves other basis states unchanged; the repeated application of $Q$, $M$ times, on the state $A|0\rangle$ gives
@@
QAE requires multiple controlled unitary gates to implement the reflection operators. We could write multiple CNOT gates (with $n$ control qubits and a single-qubit unitary) with single-qubit gates and $\leq(20n-38)$ CNOT gates [Rv2023, Eisert2000]. We consider the QAE without phase estimation that uses the maximum likelihood estimate [Suzuki2020] to integrate $\int_{0}^{c}\sin^{2}(x)$ in DQC. We compare the estimated error values against the Cramer-Rao bound [Suzuki2020], where $P(x)$ is a uniform distribution obtained by applying Hadamard gates to all qubits. The function $f(x)$ is prepared with $a=c/2^{n-1},b=c/2^{n}$ using 1 $R_{y}$ gate and $n$ controlled-Ry gates, $Ry(c/2^{n-i})$ [Suzuki2020]. It can be seen in Fig. 11 (($c=\frac{\pi}{3}$)) that DQAE can be simulated and the estimated value achieves the Cramer-Rao bound.
@@
In our simulation, we used Qiskit’s transpile function to apply a distributed multiple controlled gate and distributed unitary gates using non-local CNOT gates (Fig. 2). The DQCS can be used to obtain the resource estimation for the DQAE for various node parameters in Table 1.
@@
Table 1: Resource estimation for the distributed quantum amplitude estimation (DQAE) obtained using the DQCS where the number of nodes $\text{n}_{\text{nodes}}$, number of Grover operators $\text{n}_{\text{Grover}}$, number of entanglement generation steps $k$ for the qubit-node coupling parameter $p=(1,0.8,0.5)$, and estimation error are considered
@@
loaded to quantum state $|p\rangle=\sum_{i=0}^{2^{d-1}}\sqrt{p_{i}}|x_{i}\rangle$, where $p_{i}$ is sampled from $P(x)$ with $2^{d}$ points [Nik2020, Carrera2021]. We use the DQCS to simulate the quantum state in preparation for a probability distribution. We consider a normal distribution with $8$ qubits with the mean $\mu=1$ and $\sigma=2$ in the DQCS, with $n=1,2,4,8$ nodes, to obtain the Heilinger fidelity in Fig. 12.
Figure 12: The Heilinger fidelity of a normal distribution with eight qubits with different numbers of nodes, where $\epsilon_{d}=\epsilon_{g}=(0.002(\text{blue}),0.005(\text{orange}),0.009(\text{green}))$
@@ 5 Conclusion
We outlined a distributed quantum circuit simulator (DQCS), which uses multiple interconnected quantum processing units (QPUs) that consist of local and communication qubits. We studied foundational quantum algorithms such as the Fourier transform, quantum phase estimation, and amplitude estimation using DQCS and showed the application of dynamic quantum circuits in DQC. Recently, there has been growing interest in the realization of quantum networks and distributed quantum computing in various platforms such as superconducting qubits, NV centers [Nature2013, NV], ions [Main2025, PhysRevLett.124.110501], and neutral atoms [neutral]. The DQCS provides a platform for users to input the experiment parameters, simulate distributed quantum circuits, and obtain the estimates for quantum algorithms, including quantum phase estimation, quantum amplitude estimation, quantum state preparation of the probability distribution, etc. In the future, we aim to integrate quantum internet simulators with the DQCS to study distributed quantum algorithms in the quantum internet. We look forward to include quantum error correction and fault tolerance in the DQCS to obtain favorable polynomial scaling for the fidelity of quantum algorithms [Muralidharan2016]. We could also study distributed quantum machine learning algorithms (QML) for classification and regression [Tang2023], federated QML [Sam2021] for quantum neural networks (QNN) [Gyongyosi2019], dynamic QML for recurrent QNN’s [Gyongyosi2019].
@@ Acknowledgements
I want to thank Constantin Gonciulea, Charlee Stefanski, Vanio Markov, and David Novak for their helpful discussions, comments on the DQCS, and suggestions for improving the quality of the manuscript.